# Reasoning Language Models: A Blueprint

### v3

@@ C.3.2 Policy Update
The policy update is performed on a batch $\mathcal{D}$ of reasoning sequences. As mentioned above, the reasoning sequences can be decomposed into state-action-value triplets to then perform RL training. We distinguish between three reinforcement methods: standard RL, preference-based RL, or SFT training.
Standard Policy Gradient RL Methods. Standard policy gradient methods such as Proximal Policy Optimization (PPO) [131] or REINFORCE [1, 142] are particularly suited for tasks where trajectories are collected (online) and reliably evaluated by the q-value model $Q_{\varphi}$.
$\hat{A}(s_{t})=R_{t}+\gamma V(s_{t+1})-V(s_{t}),$
where $R_{t}$ is the immediate environment reward at step $t$, $V(s_{t})$ is the state value of of state $s_{t}$, and $\gamma$ is the discount factor. We can derive the state value easily from the q-values obtained via the q-value model or the running estimates in the MCTS as follows:
$V(s_{t+1})=\frac{1}{\gamma}Q_{\varphi}(s_{t},a_{t}),$
since rewards are sparse. The standard PPO approach trains the critic model from scratch on bootstrapped rewards for this purpose. We introduce an alternative advantage computation scheme that leverages the backpropagated values from Monte Carlo Tree Search (MCTS) in conjunction with $Q_{\varphi}$, as detailed in Algorithm 6. This integration combines MCTS’s exploration and evaluation capabilities with the RL update, enhancing robustness and efficiency in reasoning tasks.
Further regularization can be imposed on the PPO training procedure. To align the policy $\pi_{\theta}$ with a reference policy $\pi_{\text{ref}}$ (usually instantiated as $\pi_{\theta}$ before phase 2) during training, the KL divergence $\texttt{KL}(\pi_{\theta}||\pi_{\text{ref}})$ , between the two distributions can be added to the training loss. Additionally, to maintain the diversity of policy generations (and exploration during training), the entropy of the policy distribution can be enhanced by subtracting it from the loss. The entropy penalty is estimated over a batch $\mathcal{D}$ of state-action pairs $(s,a)$ where $s$ denotes a reasoning sequence and $a$ the next reasoning step. The entropy of a single completion $a$ is computed by summing the entropy of its individual tokens $a_{1:|a|}$ of $a$:
$\mathcal{L}_{H}=-\frac{1}{|\mathcal{D}|}\sum_{(s,a)\in\mathcal{D}}\sum_{a_{i}\in a}\pi_{\theta}(a_{i}|[s,a_{1:{i-1}}])\log\pi_{\theta}(a|[s,a_{1:{i-1}}]).$
Direct Preference Optimization (DPO). DPO [121] aligns the policy to user preferences expressed as pairwise comparisons between reasoning sequences. Given pairs $(s^{+},s^{-})$, where $s^{+}$ is preferred over $s^{-}$. This method may not require a process reward/value model. The loss involves the sigmoid function which we denote as $\sigma$.
Supervised Fine-Tuning (SFT). As a straightforward alternative to RL, high-value reasoning sequences can be selected to perform SFT, i.e. train the policy to maximize the likelihood of these reasoning steps. The high-value reasoning sequences may be selected as terminal nodes having the highest q-value, or highest aggregated intermediate-step values. This approach is inspired by AlphaZero-like frameworks, focusing on iteratively refining the policy to generate high-quality reasoning trajectories without requiring explicit rewards.
C.3.3 Advantage Calculation (for PPO Policy Updates)
While standard advantage computation in PPO (e.g., via Generalized Advantage Estimation (GAE) [131]) is widely applicable, we propose an alternative approach tailored to our reasoning framework in Algorithm 6. Specifically, for each state/node $s$, we leverage the q-value estimates $q(s)$ obtained during the MCTS process. They were updated in the backpropagation phase to provide a more informed estimate of the q-values incorporating the estimates of the children and potentially true reward signals from terminal paths in the tree. We expect these MCTS-derived values to be more reliable as they incorporate the ground-truth terminal reward, propagated back through the tree, ensuring that a node’s value reflects both its immediate reward and the aggregated values of subsequent child states.
@@ C.3.3 Advantage Calculation (for PPO Policy Updates)
Input: Batch $\mathcal{D}$, policy $\pi_{\theta}$, reference policy $\pi_{\text{ref}}$, learning rate $\eta$, clipping parameter $\varepsilon$, preference data $\mathcal{D}_{\text{pref}}$ for DPO.
@@
2:  Select state-action-value triplets from sequences in $\mathcal{D}$
3:  for each $(s_{t},a_{t},q_{t})\in\mathcal{D}$ do
@@
8:  Optional: add KL divergence or entropy regularization.
$\mathcal{L}_{\text{PPO}}\leftarrow\mathcal{L}_{\text{PPO}}+\lambda_{\text{KL}}\texttt{KL}(\pi_{\theta}||\pi_{\text{ref}})+\lambda_{H}\mathcal{L}_{H}.$
@@
12:  Select preference pairs of reasoning sequences in $\mathcal{D}$
13:  for each pair $(s^{+},s^{-})\in\mathcal{D}_{\text{pref}}$ do
@@
$\mathcal{L}_{\text{DPO}}=\frac{1}{|\mathcal{D}_{\text{pref}}|}\sum_{(s^{+},s^{-})}\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(s^{+})}{\pi_{\theta}(s^{-})}\right)\right).$
@@
19:  Select high-value reasoning sequences $s^{+}$ from $\mathcal{D}$
@@ C.3.4 Q-Value Model Update
During phase 2, the q-value model $Q_{\varphi}$ is also updated to track MCTS-backtracked value estimates $q^{\text{MCTS}}(s_{t})$ which should be of higher quality (thanks to the final answer verifier and score aggregation from child nodes). For each state-action pair $(s,a)$, we train the q-value model $Q_{\varphi}$ via squared error minimization, to match its q-value $Q_{\varphi}(s,a)$ as closely as possible to the corresponding MCTS-value $q^{\text{MCTS}}(s^{\prime})$ which saves the updated q-value of action $a$ taken in state $s$ leading to state $s^{\prime}$.
@@
Input: Batch $\mathcal{D}$, q-value model $Q_{\varphi}$, learning rate $\eta$.
@@
$\mathcal{L}_{q}=\frac{1}{|\mathcal{D}|}\sum_{(s,a,s^{\prime})}(Q_{\varphi}(s,a)-q^{\text{MCTS}}(s^{\prime}))^{2}$.
@@ Appendix D Data Generation
D.1 Generating Data for Phase 1 Policy Model Training
@@
For each data point, consisting of a question prompt and its associated target response comprising multiple reasoning steps $(q^{(i)},[s_{1}^{(i)},\dots,s_{n}^{(i)}])$, additional tokens are introduced to explicitly mark the boundaries of the reasoning steps. Specifically, the ’End of Intermediate Step’ (EOIS) token is defined and inserted after each reasoning step $s_{j}^{(i)}$, resulting in a modified step $s_{j}^{(i)*}$. Additionally, the ’End of Sequence’ (EOS) token is appended to the final reasoning step $s_{n}^{(i)}$, yielding $s_{n}^{(i)*}=[s_{n}^{(i)};\texttt{eos}]$. This augmentation ensures that the model can consistently identify when a final solution has been reached during inference.
@@
$y_{\text{SFT}}^{(i)}=[s_{1}^{(i)},\text{eois},\text{assistant},s_{2}^{(i)},\dots,s_{n}^{(i)},\text{eos}].$
@@
D.2 Generating Data for Phase 1 Value Model Training
@@
where $T$ is the terminal step of the (sub-) reasoning chain $\tau=(s_{i},a_{i},r_{i},s_{i+1},\ldots,s_{T},a_{T},r_{T},s_{T+1})$.
@@
Given that the input of a value model is a sequence of reasoning steps - therefore a sequence of tokens, the natural value model architecture is to use an LLM on which one adds linear layer(s) and a suitable ouput activation function. Typically, it is designed to output a scalar value $\hat{V}_{\varphi}(s_{t})\in\mathcal{C}\subseteq\mathbb{R}$.
@@ D.2.1 Binary Rewards: Modeling the Likelihood of a Correct Terminal State
To generate labels for estimating the state-value function during the training of a value model, we use MCTS with simulations till a terminal node is reached and calculate the ratio between the number of correct simulations to the number of simulations. There is one very important detail, for a trajectory $\tau=(s_{i},a_{i},r_{i},s_{i+1},\ldots,s_{T+1})$ where $s_{T+1}$ is a terminal state. By definition, the true state value function at $s_{T+1}$ is zero. However, in training the value model, we avoid instructing it to output zero for terminal states. Instead, in a supervised learning setting, we can identify terminal states and directly compare the model’s predictions against the known correct outcomes (referred to here as ”golden answers”). This comparison negates the need to rely solely on the value model to estimate the value of terminal states or to determine the reward associated with transitioning into these states. During inference, while we can still recognize terminal states, we cannot evaluate them by comparing the model’s output to a golden answer. Therefore, an alternative metric is necessary. We train the value model to predict whether transitioning to $s_{T+1}$ leads to a correct terminal outcome. By learning the relationship between a node’s content and the correctness of the resulting terminal state, the model can estimate the likelihood that a terminal state leads to a correct answer. To approximate the terminal reward during inference, we define:$r(s_{T},a_{T},s_{T+1})\approx\mathbbm{1}_{[0.5,1]}(\hat{V}_{\varphi}(s_{T+1}))$ Here $\hat{V}_{\varphi}(s_{T+1})$ represents the value predicted by the value model for the terminal state $s_{T+1}$. If this predicted likelihood exceeds a threshold (e.g., 0.5), we assign a terminal reward of 1; otherwise, we assign a reward of 0. This approach allows the value model to indirectly influence the terminal reward by predicting the likelihood of a correct outcome. Consequently, during training, terminal rewards serve as labels for terminal states in the value model. It is important to note that $\hat{V}_{\varphi}(s_{T+1})$ is not used in any other context but solely to estimate the terminal reward.
@@
$-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{V}_{\varphi}(s_{t}^{(i)}))+(1-y_{i})\log(1-\hat{V}_{\varphi}(s_{t}^{(i)}))\right]$
@@ D.2.2 Continuous and Bounded Rewards: Modeling the Expected Future Reward
Therefore, to generate labels for state-value function estimate pairs to train a value model, we use MCTS with simulations and average the outcomes of the simulations. Therefore, at each newly generated node $s$ we simulate till a terminal node is reached and we record the depth - the number of steps needed starting from $s$ (since T is not identical per trajectory). We then record the the terminal reward which in our case is $r(s_{T},a_{T})=1$ for correct and $r(s_{T},a_{T})=-1$ for incorrect answers. Discounted by the depth we can average these rewards nd obtain an estimation of the node value which serves as a label for the initial value model training.
@@
We also experimented with a tanh activation output and a linear layer with clipping of the values. However, both methods proved to be unstable in training in contrast to the scaled and shifted sigmoid layer. A tanh and sigmoid layer naturally bound the output but also push values towards the extremes, enhancing the separation between high and low value estimates. This characteristic can improve the model’s ability to distinguish between highly correct and highly incorrect states which is why we are particularly interested in these activation functions.
@@
Employing a continuous and bounded reward structure offers several benefits. Unlike binary rewards, continuous rewards provide a finer distinction between varying degrees of correctness, allowing the model to capture subtle differences in terminal states. Continuous rewards can encode more information about the quality of solutions, facilitating more informed decision-making during the search process. Bounded rewards prevent extreme values, promoting numerical stability and consistent training dynamics. However, this also shows that the choice of reward values and their scaling can significantly impact the learning process, necessitating careful calibration to ensure effective training.
@@ D.3 State Action Value Function Modeling
where $T$ denotes the terminal step of the (sub-) reasoning chain $\tau=(s_{t},a_{t},r_{t},s_{t+1},\ldots,s_{T},a_{T},r_{T},s_{T+1})$. In environments characterized by sparse rewards, where $r(s_{t},a_{t})=0$ for all $t<T$, the q-value simplifies to:

### v4

@@ C.3.2 Policy Update
The policy update, which we present in Algorithm 5, is performed on a batch $\mathcal{B}$ of reasoning sequences. As mentioned above, the reasoning sequences can be decomposed into state-action-value triplets to then perform RL training. We will discuss now three different methods (standard RL, preference-based RL, and SFT training) to improve the policy model during the second phase training.
@@
Input: Batch $\mathcal{B}$, policy $\pi_{\theta}$, reference policy $\pi_{\text{ref}}$, learning rate $\eta$, clipping parameter $\varepsilon$, preference data $\mathcal{B}_{\text{pref}}$ for DPO.
@@
2:  Select state-action-value triplets from sequences in $\mathcal{B}$
3:  for each $(s_{t},a_{t},q_{t})\in\mathcal{B}$ do
@@
8:  (Optional) Add KL divergence or entropy regularization: $\mathcal{L}_{\text{PPO}}\leftarrow\mathcal{L}_{\text{PPO}}+\lambda_{\text{KL}}\texttt{KL}(\pi_{\theta}||\pi_{\text{ref}})+\lambda_{H}\mathcal{L}_{H}$.
@@
12:  Select preference pairs of reasoning sequences in $\mathcal{B}$
13:  for each pair $(s^{+},s^{-})\in\mathcal{B}_{\text{pref}}$ do
@@
$\mathcal{L}_{\text{DPO}}=\frac{1}{|\mathcal{B}_{\text{pref}}|}\sum_{(s^{+},s^{-})}\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(s^{+})}{\pi_{\theta}(s^{-})}\right)\right).$
@@
19:  Select high-value reasoning sequences $s^{+}$ from $\mathcal{B}$
@@
Standard Policy Gradient Methods such as Proximal Policy Optimization (PPO) [135] or REINFORCE [1, 146] are particularly suited for tasks where trajectories are collected (online) and reliably evaluated by the q-value model $Q_{\varphi}$.
$\hat{A}(s_{t})=R_{t}+\gamma V(s_{t+1})-V(s_{t}),$
where $R_{t}$ is the immediate environment reward at step $t$, $V(s_{t})$ is the state value of state $s_{t}$, and $\gamma$ is the discount factor. We can derive the state value easily from the q-values obtained via the q-value model or the running estimates in the MCTS as follows:
$V(s_{t+1})=\frac{1}{\gamma}Q_{\varphi}(s_{t},a_{t}),$
since rewards are sparse. The standard PPO approach trains the critic model from scratch on bootstrapped rewards for this purpose. We introduce an alternative advantage computation scheme that leverages the backpropagated values from the MCTS in conjunction with $Q_{\varphi}$, as detailed in Algorithm 6. This integration combines MCTS’s exploration and evaluation capabilities with the RL update, enhancing robustness and efficiency in reasoning tasks.
Further regularization can be imposed on the PPO training procedure. To align the policy $\pi_{\theta}$ with a reference policy $\pi_{\text{ref}}$ (usually instantiated as $\pi_{\theta}$ before the second phase) during training, the KL divergence $\texttt{KL}(\pi_{\theta}||\pi_{\text{ref}})$ between the two distributions can be added to the training loss. Additionally, to maintain the diversity of policy generations (and exploration during training), the entropy of the policy distribution can be enhanced by subtracting it from the loss. The entropy penalty is estimated over a batch $\mathcal{B}$ of state-action pairs $(s,a)$, where $s$ denotes a reasoning sequence and $a$ the next reasoning step. The entropy of a single completion $a$ is computed by summing the entropy of its individual tokens $a_{1:|a|}$ of $a$:
$\mathcal{L}_{H}=-\frac{1}{|\mathcal{B}|}\sum_{(s,a)\in\mathcal{D}}\sum_{a_{i}\in a}\pi_{\theta}(a_{i}|[s,a_{1:{i-1}}])\log\pi_{\theta}(a|[s,a_{1:{i-1}}]).$
Direct Preference Optimization (DPO) [125] aligns the policy to user preferences expressed as pairwise comparisons between reasoning sequences. Given pairs $(s^{+},s^{-})$, where $s^{+}$ is preferred over $s^{-}$. This method may not require a process-based reward/value model. The loss involves the sigmoid function which we denote as $\sigma$.
Supervised Fine-Tuning (SFT) As a straightforward alternative to RL, high-value reasoning sequences can be selected to perform SFT, i.e. train the policy to maximize the likelihood of these reasoning steps. The high-value reasoning sequences may be selected as terminal nodes having the highest q-value, or highest aggregated intermediate-step values. This approach is inspired by AlphaZero-like frameworks, focusing on iteratively refining the policy to generate high-quality reasoning trajectories without requiring explicit rewards.
C.3.3 Advantage Calculation
While standard advantage computation in PPO (e.g., via Generalized Advantage Estimation (GAE) [135]) is widely applicable, we propose an alternative approach tailored to our reasoning framework in Algorithm 6. Specifically, for each state/node $s$, we leverage the q-value estimates $q(s)$ obtained during the MCTS process. They were updated in the backpropagation phase to provide a more informed estimate of the q-values incorporating the estimates of the children and potentially true reward signals from terminal paths in the tree. We expect these MCTS-derived values to be more reliable as they incorporate the ground truth terminal reward, propagated back through the tree, ensuring that a node’s value reflects both its immediate reward and the aggregated values of subsequent child states.
@@ C.3.4 Q-Value Model Update
During the second training phase, the q-value model $Q_{\varphi}$ is also updated (see Algorithm 7) to track the MCTS-backpropagated value estimates $q^{\text{MCTS}}(s_{t})$, which should be of higher quality (thanks to the final answer verifier and score aggregation from child nodes). For each state-action pair $(s,a)$, we train the q-value model $Q_{\varphi}$ via squared error minimization, to match its q-value $Q_{\varphi}(s,a)$ as closely as possible to the corresponding MCTS-value $q^{\text{MCTS}}(s^{\prime})$, i.e. updated q-value of action $a$ taken in state $s$ leading to state $s^{\prime}$.
@@
Input: Batch $\mathcal{B}$, q-value model $Q_{\varphi}$, learning rate $\eta$.
@@
$\mathcal{L}_{q}=\frac{1}{|\mathcal{B}|}\sum_{(s,a,s^{\prime})}(Q_{\varphi}(s,a)-q^{\text{MCTS}}(s^{\prime}))^{2}$.
@@ Appendix D Data Generation
D.1 Phase 1 Policy Model Training
@@
For each data point, consisting of a question prompt and its associated target response comprising multiple reasoning steps $(q^{(i)},[z_{1}^{(i)},\dots,z_{T+1}^{(i)}])$, additional tokens are introduced to explicitly mark the boundaries of the reasoning steps. Specifically, the ’End of Intermediate Step’ (eois) token is defined and inserted after each reasoning step $z_{j}^{(i)}$, resulting in a modified step $z_{j}^{(i)*}$. Additionally, the ’End of Sequence’ (eos) token is appended to the final reasoning step $z_{T+1}^{(i)}$, yielding $z_{T+1}^{(i)*}=[z_{T+1}^{(i)};\text{eos}]$. This augmentation ensures that the model can consistently identify when a final solution has been reached during inference.
@@
$y_{\text{SFT}}^{(i)}=[z_{1}^{(i)},\text{eois},\text{assistant},z_{2}^{(i)},\dots,z_{T+1}^{(i)},\text{eos}].$
@@
D.2 Phase 1 Value Model Training
@@
where $T$ is the terminal step of the partial reasoning chain $\tau=(s_{i},a_{i},r_{i},s_{i+1},\ldots,s_{T},a_{T},r_{T},s_{T+1})$.
@@
Given that the input of a value model is a sequence of reasoning steps - therefore a sequence of tokens, the natural value model architecture is to use a language model on which one adds linear layer(s) and a suitable output activation function. Typically, it is designed to output a scalar value $\hat{V}_{\varphi}(s_{t})\in\mathcal{C}\subseteq\mathbb{R}$.
@@ D.2.1 Binary Rewards: Modeling the Likelihood of a Correct Terminal State
To generate labels for estimating the state-value function during the training of a value model, we use MCTS with simulations till a terminal node is reached and calculate the ratio between the number of correct simulations to the number of simulations. There is one very important detail, for a trajectory $\tau=(s_{i},a_{i},r_{i},s_{i+1},\ldots,s_{T+1})$ where $s_{T+1}$ is a terminal state. By definition, the true state value function at $s_{T+1}$ is zero. However, in training the value model, we avoid instructing it to output zero for terminal states. Instead, in a supervised learning setting, we can identify terminal states and directly compare the model’s predictions against the known correct outcomes (golden answers). This comparison negates the need to rely solely on the value model to estimate the value of terminal states or to determine the reward associated with transitioning into these states. During inference, while we can still recognize terminal states, we cannot evaluate them by comparing the model’s output to a golden answer. Therefore, an alternative metric is necessary. We train the value model to predict whether transitioning to $s_{T+1}$ leads to a correct terminal outcome. By learning the relationship between a node’s content and the correctness of the resulting terminal state, the model can estimate the likelihood that a terminal state leads to a correct answer. To approximate the terminal reward during inference, we define: $r(s_{T},a_{T},s_{T+1})\approx\mathbbm{1}_{[0.5,1]}(\hat{V}_{\varphi}(s_{T+1}))$. Here $\hat{V}_{\varphi}(s_{T+1})$ represents the value predicted by the value model for the terminal state $s_{T+1}$. If this predicted likelihood exceeds a threshold (e.g., 0.5), we assign a terminal reward of 1; otherwise, we assign a reward of 0. This approach allows the value model to indirectly influence the terminal reward by predicting the likelihood of a correct outcome. Consequently, during training, terminal rewards serve as labels for terminal states in the value model. It is important to note that $\hat{V}_{\varphi}(s_{T+1})$ is not used in any other context but solely to estimate the terminal reward.
@@
$-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{V}_{\varphi}(s_{t}^{(i)}))+(1-y_{i})\log(1-\hat{V}_{\varphi}(s_{t}^{(i)}))\right],$
@@ D.2.2 Continuous and Bounded Rewards: Modeling the Expected Future Reward
To generate labels for state-value function estimate pairs to train a value model, we use MCTS with simulations and average the outcomes of the simulations. Therefore, at each newly generated node $s$ we simulate till a terminal node is reached and we record the depth - the number of steps needed starting from $s$ (since T is not identical per trajectory). We then record the terminal reward which in our case is $r(s_{T},a_{T})=1$ for correct and $r(s_{T},a_{T})=-1$ for incorrect answers. Discounted by the depth we can average these rewards and obtain an estimation of the node value which serves as a label for the initial value model training.
@@
We also experimented with a tanh activation output and a linear layer with clipping of the values. However, both methods proved to be unstable in training in contrast to the scaled and shifted sigmoid layer. A tanh and sigmoid layer naturally bounds the output but also pushes values towards the extremes, enhancing the separation between high and low value estimates. This characteristic can improve the model’s ability to distinguish between highly correct and highly incorrect states which is why we are particularly interested in these activation functions.
@@
Summary
Employing a continuous and bounded reward structure offers several benefits. Unlike binary rewards, continuous rewards provide a finer distinction between varying degrees of correctness, allowing the model to capture subtle differences in terminal states. Continuous rewards can encode more information about the quality of solutions, facilitating more informed decision-making during the search process. Bounded rewards prevent extreme values, promoting numerical stability and consistent training dynamics. However, it also shows that the choice of reward values and their scaling can significantly impact the learning process, necessitating careful calibration to ensure effective training.
@@ D.3 State Action Value Function Modeling
where $T$ denotes the terminal step of the trajectory $\tau=(s_{t},a_{t},r_{t},s_{t+1},\ldots,s_{T},a_{T},r_{T},s_{T+1})$. In environments characterized by sparse rewards, where $r(s_{t},a_{t})=0$ for all $t<T$, the q-value simplifies to: